# LiRA: Light-Robust Adversary for Model-based Reinforcement Learning in Real World

### v1

@@ 8 Conclusion
Despite these results, there are still some issues to be addressed, as discussed in Section 7. Additionally, as briefly mentioned in the main text, the disturbance-aware model can be used for MPC by introducing the domain adaptation (with the uncertainty estimator), which can contribute to reducing conservativeness. Switching between the disturbance-marginalized and disturbance-aware models depending on the accuracy of the uncertainty estimator would allow MPC to behave robustly when the estimator is not confident, and to choose the more optimal behavior when it is confident, as like the literature (Xie et al., 2022) (but more explicitly).
@@ Appendix A Auto-tuning of $\lambda$
Following the literature (Kobayashi, 2023), $\lambda$ is auto-tuned as follows:
@@ Appendix B Task configuration of worm-type robot
In the simulations, the task is performed with a worm-type robot placed in two-dimensional space in the forward and vertical directions with gravity. It is simulated by Mujoco (Todorov et al., 2012) under Gymnasium API and has a total length of 1 m and a mass of 10 kg. Its linear body is segmented according to the number of joints (i.e. six for Tail2; seven for Tail3; and eight for Tail4). The range of motion of each joint is within ±$\pm 30$ degrees and each joint is actuated by a given torque directly. Torques on the four joints located in the front (i.e. the head side) of the robot can be specified by the agent as actions. On the other hand, the torques on the 2–4 joints at the rear of the robot (i.e. the tail side) are given as disturbances. However, the maximum torque on the tail side is limited to half that of the head side. The control frequency of this robot is 100 fps, but by repeating the same torques five times, the actual frequency for the agent and adversary is given to be 20 fps (i.e. every 0.05 s). A total of 500 time steps (i.e. 25 seconds) is regarded to be an episode, and learning is completed after 200 episodes.
@@ Appendix C Task configuration of quadrupedal robot
With such a base controller design, sufficient observation data are given to RL to make MDP valid if there is no disturbance. Specifically, 84 dimensions are prepared in total: the position, angular velocity, and torque (estimated from current) of each joint; the position, velocity, and force of each leg calculated from forward kinematics and the corresponding Jacobian; the angular velocity and acceleration measured by the IMU mounted on the robot’s body; the phase of the walking trajectory based on internal time; and the previous smoothed actions. However, it was difficult to solve MPC within 20 fps on such a high-dimensional observation space, while the overlapping information is expected to be included. Therefore, as described in D, q-VAE (Kobayashi and Watanuki, 2023) compresses them to the 28-dimensional latent state space. As a result, MPC is able to consider 12 time steps ahead at 20 fps, i.e. 3 walking steps ahead.
@@ Appendix D Learning conditions
The proposed framework (and also comparisons) is implemented by PyTorch (Paszke et al., 2019). Its regularization gain to the prior is set to be $\beta=1\times 10^{-3}$ for the scalar $\lambda$ or $\beta=5\times 10^{-3}$ for the state-dependent $\lambda$. The latter is adaptively adjusted to different adversarial levels in different states and is larger than the former to allow quick disturbance suppression to match this variation. In addition, $\rho=1.5$ is set for the gap tolerance of the prediction performance in the light-robust constraint. For optimizing it, a replay buffer is utilized, as shown in Alg. 1. Its maximum capacity is 102,400 and the maximum batch size is 32. In the learning phase, all data in the buffer are replayed uniformly at random without replacement. In addition, a noise-robust stochastic gradient descent algorithm, AdaTerm (Ilboudo et al., 2023a), is employed with its default configuration for optimizing all parameters.
At first, for simplicity, the following common architecture is utilized in most models. That is, it has two fully-connected layers with 100 neurons for each. The activation function for them are the combination of Squish function (Barron, 2021; Kobayashi and Aotani, 2023) and RMSNorm (Zhang and Sennrich, 2019).
@@
By transforming it through RNF, the disturbance-aware model $p_{w}(s^{\prime},r\mid s,a;d,\theta)$ is additionally designed. This RNF has a small-sized architecture to output its parameters (for LRS flows (Dolatabadi et al., 2020)) conditioned by the disturbance: specifically, two fully-connected layers with 32 neurons for each, the activation function of which is Squaresign function (Barron, 2021; Kobayashi and Aotani, 2023). At this time, the transformability defined in RNF (Kobayashi and Aotani, 2023) is specified as 0.99, for pursuring high expressive capability. Anyway, after conditioning RNF with the disturbance, $s^{\prime}$ and $r$ generated from the disturbance-marginalized model are converted to ones from the disturbance-aware model.
@@
To auto-tune $\lambda$, which corresponds to the adversary level, the state-dependent slack variable $\Delta$ is also approximated with the common architecture. Note that its boundary is satisfied by a nonlinear transformation with Squmoid function (Barron, 2021; Kobayashi and Aotani, 2023). Note that the optimization of $\Delta$ shown in eq. (17) switches its behavior with a threshold $\epsilon=0.1$. In addition, as described in the main text, $\lambda$ is also approximated as $\lambda(s;\zeta)$ ($\zeta$ the paramters of the common architecture) in the real-robot demonstration as well as $\Delta$ in order to obtain state-dependent adversary levels.
Finally, only in the real-robot demonstration, q-VAE (Kobayashi and Watanuki, 2023) is employed to compress higher-dimensional observation into low-dimensional latent state (specifically, 28 dimensions in this paper). Specifically, an encoder with the common architecture convert the observation to the state, from which a decoder with the common architecture constructs the observation. Both the encoder and decoder actualy output the parameters of diagonal multivariate normal distributions. In q-VAE, they are optimized using $q$-logarithm (Tsallis, 1988) with $q=0.95$, which makes the state sparse and reconstructs the obseration fairly. The coefficients for all the terms to be minimized are set to be one for simplicity. Note that the state extracted by the encoder is fed into the other modules after removing its computational graph, so that biases are hardly caused by factors other than compression of the observation.
@@ Appendix E MPC conditions
The learned model is used to obtain the optimal policy with AccelMPPI (Kobayashi and Fukumoto, 2022), a type of sampling-based MPC. In each iteration, 1024 candidates for the simulations and 256 for the real-robot demonstration of action sequences up to 12 time steps ahead are sampled from the policy modeled by a diagonal multivariate normal distribution, and 1/4 of them, which are expected to perform well, are selected by rejecting others. The selected candidates are evaluated by predicting the future states and rewards using the model. According to the predicted rewards, the policy is updated to proceed to the next iteration. Although the details of its update law is omitted in this paper (see the literature (Kobayashi and Fukumoto, 2022)), the step size is 0.25, the inverse temperature is 1, the update ratio of the policy to generate negative results is 1, and the slowdown gain to adjust a accelerated gradient is 0.5.

### v2

@@ 8 Conclusion
Despite these results, there are still some issues to be addressed, as discussed in Section 7. Additionally, as briefly mentioned in the main text, the disturbance-aware model can be used for MPC by introducing the domain adaptation (with the uncertainty estimator), which can contribute to reducing conservativeness. Switching between the disturbance-marginalized and disturbance-aware models depending on the accuracy of the uncertainty estimator would allow MPC to behave robustly when the estimator is not confident, and to choose the more optimal behavior when it is confident, as like the literature [68] (but more explicitly). Alternatively, in the direction of improving robustness, the uncertainty of disturbance-marginalized model can be explicitly taken into account by appropriately combining a robust MPC [43, 44]. However, as discussed in the main text, excessive robustness makes the policy conservative, so it is desirable to adaptively determine how much uncertainty to take into account like the literature [69]. After these further improvements, a comprehensive investigation with existing robust methods will be conducted to show the practical value of LiRA clearly.
@@ Appendix A Auto-tuning of $\lambda$
Following the literature [15], $\lambda$ is auto-tuned as follows:
@@ Appendix B Task configuration of worm-type robot
In the simulations, the task is performed with a worm-type robot placed in two-dimensional space in the forward and vertical directions with gravity. It is simulated by Mujoco [16] under Gymnasium API and has a total length of 1 m and a mass of 10 kg. Its linear body is segmented according to the number of joints (i.e. six for Tail2; seven for Tail3; and eight for Tail4). The range of motion of each joint is within ±$\pm 30$ degrees and each joint is actuated by a given torque directly. Torques on the four joints located in the front (i.e. the head side) of the robot can be specified by the agent as actions. On the other hand, the torques on the 2–4 joints at the rear of the robot (i.e. the tail side) are given as disturbances. However, the maximum torque on the tail side is limited to half that of the head side. The control frequency of this robot is 100 fps, but by repeating the same torques five times, the actual frequency for the agent and adversary is given to be 20 fps (i.e. every 0.05 s). A total of 500 time steps (i.e. 25 seconds) is regarded to be an episode, and learning is completed after 200 episodes.
@@ Appendix C Task configuration of quadrupedal robot
With such a base controller design, sufficient observation data are given to RL to make MDP valid if there is no disturbance. Specifically, 84 dimensions are prepared in total: the position, angular velocity, and torque (estimated from current) of each joint; the position, velocity, and force of each leg calculated from forward kinematics and the corresponding Jacobian; the angular velocity and acceleration measured by the IMU mounted on the robot’s body; the phase of the walking trajectory based on internal time; and the previous smoothed actions. However, it was difficult to solve MPC within 20 fps on such a high-dimensional observation space, while the overlapping information is expected to be included. Therefore, as described in D, q-VAE [41] compresses them to the 28-dimensional latent state space. As a result, MPC is able to consider 12 time steps ahead at 20 fps, i.e. 3 walking steps ahead.
@@ Appendix D Learning conditions
The proposed framework (and also comparisons) is implemented by PyTorch [70]. Its regularization gain to the prior is set to be $\beta=1\times 10^{-3}$ for the scalar $\lambda$ or $\beta=5\times 10^{-3}$ for the state-dependent $\lambda$. The latter is adaptively adjusted to different adversarial levels in different states and is larger than the former to allow quick disturbance suppression to match this variation. In addition, $\rho=1.5$ is set for the gap tolerance of the prediction performance in the light-robust constraint. For optimizing it, a replay buffer is utilized, as shown in Alg. 1. Its maximum capacity is 102,400 and the maximum batch size is 32. In the learning phase, all data in the buffer are replayed uniformly at random without replacement. In addition, a noise-robust stochastic gradient descent algorithm, AdaTerm [71], is employed with its default configuration for optimizing all parameters. AdaTerm is expected to automatically exclude outliers hidden in the worst-case data that are likely to be affected by MMB at training time.
At first, for simplicity, the following common architecture is utilized in most models. That is, it has two fully-connected layers with 100 neurons for each. The activation function for them are the combination of Squish function [72, 46] and RMSNorm [73].
@@
By transforming it through RNF, the disturbance-aware model $p_{w}(s^{\prime},r\mid s,a;d,\theta)$ is additionally designed. This RNF has a small-sized architecture to output its parameters (for LRS flows [74]) conditioned by the disturbance: specifically, two fully-connected layers with 32 neurons for each, the activation function of which is Squaresign function [72, 46]. At this time, the transformability defined in RNF [46] is specified as 0.99, for pursuring high expressive capability. Anyway, after conditioning RNF with the disturbance, $s^{\prime}$ and $r$ generated from the disturbance-marginalized model are converted to ones from the disturbance-aware model.
@@
To auto-tune $\lambda$, which corresponds to the adversary level, the state-dependent slack variable $\Delta$ is also approximated with the common architecture. Note that its boundary is satisfied by a nonlinear transformation with Squmoid function [72, 46]. Note that the optimization of $\Delta$ shown in eq. (18) switches its behavior with a threshold $\epsilon=0.1$. In addition, as described in the main text, $\lambda$ is also approximated as $\lambda(s;\zeta)$ ($\zeta$ the paramters of the common architecture) in the real-robot demonstration as well as $\Delta$ in order to obtain state-dependent adversary levels.
Finally, only in the real-robot demonstration, q-VAE [41] is employed to compress higher-dimensional observation into low-dimensional latent state (specifically, 28 dimensions in this paper). Specifically, an encoder with the common architecture convert the observation to the state, from which a decoder with the common architecture constructs the observation. Both the encoder and decoder actualy output the parameters of diagonal multivariate normal distributions. In q-VAE, they are optimized using $q$-logarithm [75] with $q=0.95$, which makes the state sparse and reconstructs the obseration fairly. The coefficients for all the terms to be minimized are set to be one for simplicity. Note that the state extracted by the encoder is fed into the other modules after removing its computational graph, so that biases are hardly caused by factors other than compression of the observation.
@@ Appendix E MPC conditions
The learned model is used to obtain the optimal policy with AccelMPPI [39], a type of sampling-based MPC. In each iteration, 1024 candidates for the simulations and 256 for the real-robot demonstration of action sequences up to 12 time steps ahead are sampled from the policy modeled by a diagonal multivariate normal distribution, and 1/4 of them, which are expected to perform well, are selected by rejecting others. The selected candidates are evaluated by predicting the future states and rewards using the model. According to the predicted rewards, the policy is updated to proceed to the next iteration. Although the details of its update law is omitted in this paper (see the literature [39]), the step size is 0.25, the inverse temperature is 1, the update ratio of the policy to generate negative results is 1, and the slowdown gain to adjust a accelerated gradient is 0.5.